OpenStudy (anonymous):

the following is the graph of an equation 4y2 -4x=16 find the x and y intercepts

OpenStudy (anonymous):

Is that "4y squared", like below? \[4y ^{2}\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

x intercept is where the curve crosses the x axis, so it's where y=0. Sub in y=0 to the curve equation: \[4y ^{2} - 4x = 16\] 4(0) - 4x = 16 -4x = 16 ----->>>>> x = -4 So the x intercept is (-4, 0) or just -4, depending on how they want you to respond. Do the same for the y intercept, where x = 0. \[4y ^{2} - 4(0) = 16\] \[4y ^{2} = 16\] \[y ^{2} = 4\] Then y = 2.... AND negative 2, since both satisfy the equation. Y intercepts exist at (0,2) and (0,-2)

OpenStudy (anonymous):

thanks so much !!! =)

OpenStudy (anonymous):

Also, notice they are messing with you by turning the graph sideways so that the y axis is horizontal... it doesn't change the problem any, so turn your book if you like it with y vertical :)

OpenStudy (anonymous):

your right lol thanks again =)

OpenStudy (anonymous):

Do you have to close the question when you're done? I'm new here...

OpenStudy (anonymous):

im new too lol just signed up

OpenStudy (anonymous):

well, nice to know it isn't just me. Are you able to click the thing at the top of this that says "The question is open" ? I'm guessing only the person who asked can close it. No big deal... I guess I should have read the instructions after I signed up :)

OpenStudy (anonymous):

lol same here i actually closed the question just now so that i may ask another one

OpenStudy (anonymous):